please help solve fractions.

-2x+ one-third y +4=0; x=2

Answer by rfer(16322)

-2x+1/3y+4=0
-2(2)+1/3y+4=0
-4+1/3y+4=0
1/3y=0
multiply both sides by 3
1/3y*3=0*3
y=0